1. Does the discrepancy between net income and operating cash flows grow larger or smaller?

2. What are the three largest adjustments to net income, other than depreciation?

3. Did the company make any new investments in the past year? Describe them.

4. Did the company pay any dividends? If so, how much?

5. Did the company buy back any shares of its own? If so, how many shares? What is the total dollar value?

Average Char 2016 vs. 2017 44% 34% 66% 10% 133% 14% 105% 396% 277% 463% 2015 2016 2017 2015 vs. 2016 26,232 16529.00 9703.00 2930.00 6773.00 1234.00 9.00 5530.00 5460.00 10,990.00 22% 16% 34% 2% 72% 3% 132% 203% 109% 332% 18,338 526,487 S Net Sales/Sales Revenue Cost of Goods Sold Gross Profit Selling, General&Administrative Operating Income Interest Expenses Other Expenses Earnings bf Taxes Tax Expense Net Income 16901.00 9586.00 3444.00 6142.00 1134.00 15.00 5023.00 1381.00 2% 1% 15% 10% 9% 160% 10% 495% 202% 12577.00 5761.00 3122.00 2639.00 1321.00 305.00 1013.00 366.00 647.003,642.00 $Kraft Heinz Operating Cash Flows Investing Cash Flows Financing Cash Flows 2015 2,467.00 S 2016 5,238.00 $ 1.113.00) 4,621.00) $ 2017 2015 vs 2016 2016 vs 2017 527.00 1,156.00 4,226.00 112.32% 88.53% 145.38% -89.94% 203.86% -8.55% (9,704.00) 10,183.00$
Answer #1

1 Simple way to find out whether discrepancy between net income and operating Cash flow increased or decreased, is to compare it in relative terms as follows

2015 :- net income / operating Cash flow = 647/2467 = 26.23%

2016 :- net income / operating Cash flow = 3642/5238 = 69.53%

2017 :- 10990/527 = 2085%.

Note that as this ratio will be more closer to 100% discepency will be lower .

Here we can see that from 2015 to 2016 this discrepancy get smaller but fin 2017 this, discrepancy again grow larger.

2. Three largest adjustment to Net income other than depreciation are :-

A . Any gain or loss on sale of any asset or investment .

B. Any increase or decrease of current assets.

C. Any increase or decrease of current liabilities .

3 . Yes , company made new investment in the year 2015 and 2016 as investing Cash flow was negative in 2015 as well as in 2016 which shows that Cash have been used for purchasing new investment.

4 . With he available information, it cannot be said that whether company has paid any dividend and the amount of such dividend . We can see that financing cash flow is negative which is an indicator of payment of dividend . But it can also be negative if any debt have been repaid or if we have purchssed our own stock. Analysis of balance sheet is also required for this purpose.

5. Again it cannot be said with certainty that whether company has buyback it's share on the basis of the given information. Although financing Cash flow is negative and this is an indicator of buyback of share but it is also an indicator of payment of dividend or repayment of debt . Analysis of balancehseet is required in such a case.